OpenStudy (anonymous):

9p+8=10p+7 Is This Equation True,False Or Open. Please Explain Why. :)

jimthompson5910 (jim_thompson5910):

when you solve this, what do you get

OpenStudy (anonymous):

Im Not So Sure On How To Solve It

OpenStudy (anonymous):

Or The First Step

jimthompson5910 (jim_thompson5910):

the goal is to isolate P

OpenStudy (anonymous):

Right

jimthompson5910 (jim_thompson5910):

so move everything that has a p to one side, everything else to the other

OpenStudy (anonymous):

Okay So 9p and 10p

OpenStudy (anonymous):

Would I Add Those Two Togather

OpenStudy (anonymous):

9P+8=10P+7 9P-10P=7-8 -1P=-1 P=1

OpenStudy (anonymous):

you have to isolate p and find its value

OpenStudy (anonymous):

Okay Well Thanks I Know How To Do It But Sometimes Im Just Confused Thank You Very Much @jim_thompson5910 and @ArbabShah

OpenStudy (anonymous):

anytime thats what we are here for :)

OpenStudy (anonymous):

So Then That Equation Would Be True Then Right

jimthompson5910 (jim_thompson5910):

is it always true? for every value of p?

OpenStudy (anonymous):

I have No Idea What Ur Asking Sorry

jimthompson5910 (jim_thompson5910):

I recommend you going back to the lesson to learn how to solve equations and look up the terms (like what it means when an equation is true)

OpenStudy (anonymous):

@Aminah14 yes

OpenStudy (anonymous):

Okie Sounds Like A Plan @jim_thompson5910

OpenStudy (anonymous):

And Thank You @ArbabShah